A person's blood glucose level and diabetes are closely related. Let x be a random variable measured in milligrams of glucose per deciliter (1/10 of a liter) of blood. Suppose that after a 12-hour fast, the random variable x will have a distribution that is approximately normal with mean μ = 88 and standard deviation σ = 23. Note: After 50 years of age, both the mean and standard deviation tend to increase. For an adult (under 50) after a 12-hour fast, find the following probabilities. (Round your answers to four decimal places.)

(a) x is more than 60


(b) x is less than 110


(c) x is between 60 and 110


(d) x is greater than 140 (borderline diabetes starts at 140)
Answer #1

Solution :

Given that ,

mean = \mu = 88

standard deviation = \sigma = 23

a)

P(x > 60) = 1 - P(x < 60)

= 1 - P((x - \mu ) / \sigma < (60 - 88) / 23)

= 1 - P(z < -1.22)

= 1 - 0.1112 Using standard normal table.

Probability = 0.8888

b)

P(x < 110) = P((x - \mu ) / \sigma < (110 - 88) / 23)

= P(z < 0.96)

= 0.8315 Using standard normal table,

Probability = 0.8315

c)

P(60 < x < 110) = P((60 - 88)/ 23) < (x - \mu ) /\sigma  < (110 - 88) / 23) )

= P(-1.22 < z < 0.96)

= P(z < 0.96) - P(z < -1.22)

= 0.8315 - 0.1112 Using standard normal table,  

Probability = 0.7203

d)

P(x > 140) = 1 - P(x < 140)

= 1 - P((x - \mu ) / \sigma < (140 - 88) / 23)

= 1 - P(z < 2.26)

= 1 - 0.9881 Using standard normal table.

Probability = 0.0119